OpenStudy (tdskitz):

Solve -1 over 2 x + 2 = -x + 7 A) 10 B) 5/2 (5 over 2) C) 2/5 (2 over 5) D) 20

OpenStudy (lime):

\[\frac{ -1 }{ 2x }+ 2 = -x + 7\] Does your problem looks like this?

OpenStudy (tdskitz):

no

OpenStudy (tdskitz):

\[-\frac{ 1 }{ 2 } x + 2 = -x + 7\]

OpenStudy (madisongoode):

Are you trying to find out what "x" is?

OpenStudy (tdskitz):

I don't know that's why I'm here x.x

OpenStudy (madisongoode):

Well, I know "x" equals 10. But I don't know if that's what they want...

OpenStudy (tdskitz):

I dont know if this will help but its from Algebra 1 FLVS 2.04

OpenStudy (madisongoode):

Okay.

OpenStudy (ace-n-it):

A

OpenStudy (tdskitz):

alright ill go with a, but i have 3 more

OpenStudy (lime):

Yeah i got a also.

OpenStudy (madisongoode):

Same.